OpenStudy (anonymous):

Help please, Use properties of logarithms to rewrite the formula E=log I^2/MR in terms of logarithms that do not contain fractions or powers.

OpenStudy (anonymous):

I think the equation you wrote is \[E = \log(\frac{I^2}{M*R})\]If it is, then it is equivalent to \[E = 2\log(I) - \log(M) - \log(R)\]

OpenStudy (anonymous):

Thanks, I thought it would be E=2log(I)-log(M)+log (R)